OpenStudy (anonymous):

Show that (sin 45)squared+(cos 45)squared =1

OpenStudy (anonymous):

We know sin(45 = sqrt(2)/2 and cos(45 = sqrt(2)/2

OpenStudy (anonymous):

Okay. Oooooohhhhh. I understand now. Thanks so muh

OpenStudy (anonymous):

\[\frac{ \sqrt(2) }{ 2 }^{2}= \frac{ 1 }{2 }\] (sqrt(2)/2)^2 + (sqrt(2)/2)^2 = 1

OpenStudy (anonymous):

Also remember that this is a pythagorean identity cos^2 + sin^2 = 1